Fermi National Laboratory
 

Operation Summary for April 16, 2010 - April 19, 2010

 

The Integrated Luminosity for the period from 4/5/10 to 4/12/10 was 49.33 pb-1.
NuMI reported receiving 6.1E18 protons on target during this same period.

Friday April 16

The day shift began with Operations monitoring store 7746 (178.83E30), the Recycler stash (228.8E10), and the Antiproton Source stack (6.9E10), with the Linac, Booster, and Main Injector operating normally, and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, repairs to the Linac Buncher and calibration of the Main Injector RF 3-way valve accounted for 2 hours of downtime.

During the evening shift, Operations terminated store 7746 at 7:04 PM and put the TeV into shot setup at 7:15 PM. Operations established store 7747 at 8:21 PM with an AIL of 402.4E30 (A new record). An LCW water leak was found and repaired in the AP-30 stub, this caused 2 hours of stacking downtime for Pbar.

Saturday April 17

The midnight shift began with Operations monitoring store 7747 (191.88E30) and the stash (50.9E10), with Pbar stacking (22.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a vacuum pump faulted off in the 8 Gev line causing minimal downtime; pump was reset.

During the day shift, Operations monitored store 7747.

During the evening shift, Operations terminated store 7747 at 4:41 PM and put the TeV into shot setup and established store 7748 at 6:27 PM with an AIL of 373E30.

Sunday April 18

The midnight shift began with Operations monitoring store 7748 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, the MI-52 LCW pumps tripped off at 6:32 AM. Operations prepared the Main Injector for access (dumping the RR stash of 311E10 and turning beam off to Pbar, NuMI and MiniBooNE) to search for and repair the water leak.

During the day shift, Operations monitored store 7748. The Main Injector water leak was found and repaired by 11:37 AM. At 3:18 PM all Main Injector LCW pumps were back on. Operations noticed that vacuum was bad in the P150 line, since this is in the same area as the water leak, vacuum equipment got wet.

During the evening shift, experts were called in, to find and repair the vacuum problem in the P150 line. A vacuum leak was found and repaired by 11:34 PM.

Monday April 19

The midnight shift began with Operations monitoring store 7748.

During the midnight shift, Operations established beam to NuMI at 1:51 AM, beam to MiniBooNE was delayed due to a power supply problem, but was established at 3:17 AM. Once the pump-down was completed in the P150 line, beam was established to Pbar at 4:15 AM. Tuning on the machines continued through the rest of the shift. Near the end of the shift, the Ecool transfer line solenoid tripped off, Operations could not get it back on and contacted experts.

The Plans for Monday

The plan for today is to repair the Ecool solenoid and to stack, stash and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/19/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory